OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) Solve the quantity x minus 5, end quantity, over x, minus 1, equals the quantity x plus 3, end quantity, over x. 8 0 −1 −8 Question 2 (Multiple Choice Worth 1 points) Solve 2 over 8 times x squared equals 1 over 8 times x. −2 −1 1 2 Question 3 (Multiple Choice Worth 1 points) Solve the quantity x plus 4, end quantity, over x, plus 1, equals the quantity x plus 6, end quantity, over x. 0 2 4 6 Question 4 (Multiple Choice Worth 1 points) Solve 1 over 6 equals 3 over the quantity x plus 12, end quantity. 6 8 12 30

OpenStudy (anonymous):

Part 1: Explain how you calculate the restricted value(s) of x for the equation the fraction 3 over the binomial x plus 2 plus the fraction 1 over 2x equals the fraction 4 over the binomial x plus 2 and what those value(s) represent. Part 2: Explain which method (cross-multiplication or LCD) you would use to solve the equation and why. Part 3: Using complete sentences, explain the steps to solve the equation using the method you chose in step 2. Provide the solution to the equation. Can anyone please help me? I want to check my answers, but I don't get this question right here^^. My answers are B,D,B, and A.

OpenStudy (anonymous):

Oops sorry I meant C,D,B, and A for my answers :)

OpenStudy (kira_yamato):

[(x - 5)/x] - 1 = (x + 3)/x (x - 5) - x = x + 3 x = -8

OpenStudy (kira_yamato):

Can you give me the 2nd one in eqn form? Thx

OpenStudy (anonymous):

Thanks, and sure! :D

OpenStudy (anonymous):

\[2/8x^2=1/8x\]

OpenStudy (kira_yamato):

Is the x^2 with the 8?

OpenStudy (anonymous):

yes it is

OpenStudy (kira_yamato):

And what about the x?

OpenStudy (anonymous):

Also with the 8

OpenStudy (kira_yamato):

Ok So 2/(8x^2) = 1/(8x) Multiplying throughout by 8x^2 2 = x

OpenStudy (kira_yamato):

[(x + 4)/x] + 1 = (x + 6)/x x + 4 + x = x + 6 2x - x = 6 - 4 x = 2

OpenStudy (anonymous):

Ok, thanks! The next one would be \[x+4/x +1= x+6/x\]

OpenStudy (kira_yamato):

1/6 = 3/(x + 12) Cross multiply: x + 12 = 18 x = 6

OpenStudy (anonymous):

Thank you so much!

OpenStudy (kira_yamato):

No problem